# GUIDE: Enhancing Gradient Inversion Attacks in Federated Learning with Denoising Models

### v1

@@ Abstract
keywords:
Federated Learning , Gradient Inversion Attacks , Reconstruction Attacks , Privacy Attacks , Privacy-Preserving Machine Learning
[…]
[unisa]organization=Department of Information and Electrical Engineering and Applied Mathematics, University of Salerno, addressline=Via Giovanni Paolo II, 132, city=Fisciano, postcode=84084, state=SA, country=Italy
1 Introduction
@@
Federated Learning (FL) McMahan et al. (2017) offers an alternative to traditional centralized Machine Learning (ML) by enabling multiple participants to collaboratively train shared models without exchanging the training data Bellavista et al. (2021). In an FL setting, a central server distributes a global model to participating clients that locally compute updates (i.e., gradients or weights difference) using their private dataset. The updates are then sent to the server, which aggregates them to obtain an updated version of the global model that is sent back to the clients. This process is repeated until convergence.
@@
Although FL enhances privacy protection compared to centralized approaches as clients’ data are kept on-site Mothukuri et al. (2021), it is still vulnerable to several security and privacy threats Rodríguez-Barroso et al. (2023); Kumar et al. (2024); Yin et al. (2021b). In particular, privacy attacks, which are the main focus of this work, aim to infer sensitive information about the clients, such as training data or their properties, by exploiting the intermediate updates sent to the server Kumar et al. (2024); Yin et al. (2021b); Carletti et al. (2025); Du et al. (2025).
@@
In FL scenarios, server-client communications can be protected through end-to-end encryption, making attacks by external adversaries more difficult. However, this does not prevent an honest-but-curious server from breaking clients’ privacy without being detected, as it follows the regular training protocol while seeking to gain additional information Kumar et al. (2024); Carletti et al. (2025). The server has access to the global model and client updates. Additionally, it is capable of executing computationally intensive reconstruction methods.
@@
Within this threat model, Gradient Inversion Attacks (GIAs) aim to reconstruct the private client data used to compute the model update Carletti et al. (2025). These attacks rely on optimization techniques to reconstruct client’s data, with the objective to estimate input values that produce updates closely matching those shared by the victim client Zhu et al. (2019); Geiping et al. (2020). Optimization strategies may also employ some pre-trained generative models like Generative Adversarial Networks (GANs) or diffusion models as strong priors to constrain the optimization process Jeon et al. (2021); Li et al. (2022); Gu et al. (2024). In the latter case, the attacker must have access to a surrogate dataset to train these models. These datasets should approximate target client’s data distribution (e.g., if the learning task consist of recognizing faces, then the attacker should have some public photos of other individuals). This assumption is widely accepted in the literature Li et al. (2022); Gu et al. (2024); Yue et al. (2023); Fang et al. (2023); Yang et al. (2023a); Xue et al. (2023); Yao et al. (2025); Carletti et al. (2025).
@@
Figure 1: Comparison of reconstruction results on images when the model update is calculated over 256 images at $224\times 224$ resolution using the VGG-16 model. (a) Reconstruction with CPA Kariyappa et al. (2023), (b) Reconstruction with CPA Kariyappa et al. (2023) enhanced with GUIDE (ours), (c) Original Images.
GIAs have been demonstrated to be more effective when model updates are computed over small batch sizes or when the data is low-dimensional Zhu et al. (2019); Yin et al. (2021a); Geiping et al. (2020); Li et al. (2022); Yang et al. (2023a); Gu et al. (2024). However, FL clients usually compute gradients over large batch sizes, and data is high-dimensional (e.g., high-resolution images), making data reconstruction even more challenging. As shown in Figure 1a, when larger batches of high-resolution images are used, existing GIAs often generate noisy approximations of the client’s original private images. This limitation can be overcome if the attacker has access to a surrogate dataset to train tailored ML models aiming to enhance reconstruction quality. To this purpose, recent advancements in generative neural networks, particularly diffusion models Yang et al. (2023b), hold significant promise to make GIAs more effective.
@@
We implemented GUIDE and conducted extensive evaluations across various state-of-the-art attacks, using different FL algorithms, models, and datasets, while also considering defensive mechanisms. Furthermore, we recognize that commonly used perceptual metrics for assessing GIAs may not be entirely suitable for evaluating privacy leakage in reconstructed images because they do not fully align with human perception Sun et al. (2024). To address this issue, we propose using DreamSim Fu et al. (2023), which has recently been established as the state-of-the-art perceptual similarity metric Ghazanfari et al. (2024). Experimental results demonstrate that GUIDE significantly enhances the quality of reconstructed data in GIAs, generating images that are both more realistic and semantically closer to the original ones. Specifically, it achieves up to 46% higher perceptual similarity according to the DreamSim metric, with even further gains when certain defensive strategies are applied. Figure 1b visually illustrates how GUIDE enhances the effectiveness of the Cocktail Party Attack (CPA) Kariyappa et al. (2023). These improvements, achieved under the same threat model, directly correlate with increased privacy leakage, even when updates are computed over large batch sizes and high-dimensional data.
@@
  • We demonstrate that the DreamSim Fu et al. (2023) metric better assess the privacy leakage of reconstructed images, offering a more perceptually aligned evaluation of GIAs in FL;
@@
Organization. The remainder of the paper is organized as follows. Section 2 provides the background and review of related work on GIAs. Section 3 introduces the considered threat model. Section 4 presents GUIDE, whose effectiveness is extensively evaluated in Section 5. Finally, Section 6 draws our conclusions. Notations and symbols used throughout the paper are clarified in Table 1.
@@
Table 1: Formal notations used in the paper.
2 Background & Related Work
@@ 2.1 Federated Learning
The primary algorithms used in FL are Federated Stochastic Gradient Descent (FedSGD) McMahan et al. (2017) and Federated Averaging (FedAVG) McMahan et al. (2017), which mainly differ in how the selected clients $\mathcal{K}\subseteq C$ compute and transmit updates at each iteration to train the global model. In FedSGD McMahan et al. (2017), each client performs a single local epoch ($E=1$) by sampling a batch from its private dataset ($B\leq N_{c}$) and computing the aggregated gradient $\nabla_{c}=\sum_{i=1}^{B}\nabla\mathcal{L}(x_{c,i},y_{c,i};\theta_{t})$. This computed gradient is then sent to the server, which aggregates the gradients from all selected clients to update the global model as follows:
@@
where $\eta$ is the global learning rate. While FedSGD enables frequent global model updates, it incurs significant communication overhead due to the continuous exchange of gradients McMahan et al. (2017). FedAVG McMahan et al. (2017) improves communication efficiency by allowing clients to perform multiple local training steps ($E>1$) before sharing updates with the server. In each communication round $t$, a client $c\in\mathcal{K}$ initializes its local model as $\theta_{c}^{(t,0)}=\theta_{t}$ and performs $\tau=E\cdot N_{c}/B$ gradient descent steps using a learning rate $\eta$, updating its local model to $\theta_{c}^{(t,\tau)}$. Once local training is completed, the client computes the model update as:
@@
2.2 Gradient Inversion Attacks
@@
where $\mathcal{L}_{\text{grad}}$ denotes a loss function that quantifies the discrepancy between the shared update $\nabla$ and the dummy update $\nabla^{\prime}$ computed with respect to the reconstructed input $X^{\prime}$ and labels $Y^{\prime}$. $\mathcal{R}_{\text{aux}}$ is a regularization term that enforces specific constraints on the reconstructed input $X^{\prime}$. The optimization variables $X^{\prime}$ and $Y^{\prime}$ can be initialized either randomly or leveraging prior knowledge relevant to the problem Carletti et al. (2025).
One of the first attacks is Deep Leakage from Gradients (DLG) Zhu et al. (2019), which targets small batches and simple architectures under the FedSGD protocol. Subsequent works have demonstrated the effectiveness of similar methods in the presence of more complex FL schemes, such as FedAVG Dimitrov et al. (2022); Yue et al. (2023). Progress in this domain has generally involved refining the optimization pipeline with additional priors and constraints. For instance, InvGrad Geiping et al. (2020) uses a combination of cosine distance and total variation regularization, thereby enabling the inversion of deeper networks and larger batches. GradInversion Yin et al. (2021a) further exploits Batch Normalization (BN) statistics to guide reconstructions, aligning the activations produced by candidate inputs with the global BN statistics stored in the model. Meanwhile, some works isolate label reconstruction as a separate challenge, which can be addressed analytically (e.g., assuming unique labels per batch). In another direction, CPA Kariyappa et al. (2023) casts gradient inversion in Fully Connected (FC) layers as a blind source separation problem, first recovering private embeddings by unmixing the FC-layer gradients via techniques akin to independent component analysis, and then reconstructing images through an auxiliary optimization.
Some other attacks Jeon et al. (2021); Li et al. (2022); Fang et al. (2023); Xu et al. (2023) leverage pre-trained generative models to reduce the complexity of searching directly in the high-dimensional image space. Concretely, these approaches restrict the optimization process to a latent code $z^{\prime}\in\mathcal{Z}$, under the assumption of having a generator $G\colon\mathcal{Z}\to\mathcal{X}$ that maps latent vectors to candidate images. The server-side updates $\nabla^{\prime}$ are then computed by substituting $G(z^{\prime})$ into the client’s model in place of the actual private inputs. Formally, one solves:
@@
Recently, Gu et al. (2024) have proposed a novel GIA based on the latent-space optimization principle (Eqn. 7), employing a diffusion model as the pre-trained generative model instead of a GAN. It is worth noting that the diffusion model is used as a generative prior, guiding the iterative synthesis process from pure noise by incorporating the shared gradient as a directive condition at each sampling step. This significantly differs from our methodology that leverages the diffusion model as a denoising tool specialized for the underlying GIA. Additionally, Gu et al. (2024) was validated on model updates computed from a single image (a batch size of one), which is an unrealistic setting for practical FL deployments.
2.3 Augmented Optimization Attacks
Other strategies seek to improve GIAs by incorporating image restoration modules, particularly when defensive measures based on gradient obfuscation are in place Carletti et al. (2025). Highly-Compressed Gradient Leakage Attack (HC-GLA) Yang et al. (2023a) combines standard gradient-based reconstruction steps with a specialized Convolutional Neural Network (CNN) for noise suppression. This method differs from state-of-the-art image restoration models, as it is a CNN trained on a limited dataset of only a few thousand images, which restricts its overall potential and effectiveness in comparison to more advanced approaches. Moreover, it is primarily evaluated on single-image batches and may struggle to scale or generalize to diverse data distributions or larger batch sizes.
Reconstructing from Obfuscated Gradient (ROG) Yue et al. (2023) attack introduces a two-step GIA designed to enhance both convergence and visual fidelity. First, the attack operates in a reduced-dimensional space by encoding the input data, simplifying the optimization objective, and accelerating convergence. Once a suitable representation is found, a GAN is used to restore the optimized embedding into the original high-dimensional image space. Unlike GUIDE, this GAN is not specifically trained on the intermediate noisy reconstructions produced by the initial attack step. This limits its adaptability and results in a significant performance drop when targeting deeper neural architectures.
@@ 2.3 Augmented Optimization Attacks
3 Threat Model
In this section, we provide a comprehensive definition of the threat model adopted in our work, specifying the nature of the adversary, along with its objectives, knowledge, and capabilities. In line with the literature on GIAs Carletti et al. (2025); Du et al. (2025), we consider the central server as the adversary.
Adversary Goal. The main goal of the central server is to break clients’ privacy, by gaining information about their private data used for the training.
@@ 3 Threat Model
In some prior works, the use of a surrogate dataset is not explicitly stated; however, this assumption is implicitly made through the application of pre-trained generative models. Specifically, this threat model closely align with one of the threat models defined in Carletti et al. (2025), followed by several works in literature Li et al. (2022); Gu et al. (2024); Yue et al. (2023); Fang et al. (2023); Yang et al. (2023a); Xue et al. (2023); Yao et al. (2025).
Adversary Capability. The central server in FL could be either active malicious or honest-but-curious Carletti et al. (2025). An active malicious server intentionally deviates from the protocol to compromise client privacy by manipulating the global model to facilitate data reconstruction. In contrast, we adopt the honest-but-curious server, a more realistic threat model in FL deployments. This reflects scenarios where the server faithfully follows the FL protocol, but passively attempts to infer private information from clients’ updates Carletti et al. (2025). It is worth noting that, under this threat model, clients are unable to detect the server’s malicious behavior.
4 GUIDE
@@
4.1 Intuition behind GUIDE
Several works have experimentally shown the effectiveness of GIAs to reconstruct an approximated version of client’s input, especially when additional priors constrain the optimization procedure Carletti et al. (2025). However, there is no theoretical guarantee that these reconstructions correspond to the data used by the client during the local training. Indeed, GIAs based on gradient-matching optimization procedures (e.g., those based on Equations 6 and 7) are inherently ill-posed mathematical problems since they potentially admit several solutions consistent with the observed client’s gradient Zhu and Blaschko (2021).
The reconstructed inputs obtained through GIAs often exhibit noisy artifacts, which can be attributed to the fact that the optimization may place the reconstructed points in lower-density regions of the data manifold, where the original client data is less likely to reside. Thus, the recovered points may satisfy the observed gradients but still deviate from the true data distribution. Given these considerations, we leverage generative image enhancement models trained with a surrogate dataset to move approximated input reconstructions into regions of the data space that are more likely to contain the original client inputs Vincent (2011). This “guides” the optimization process toward high-density areas of the learned distribution (hence the name of the method), reducing ambiguity in the reconstruction and improving the overall fidelity of the recovered samples.
4.2 GUIDE Setup

### v2

@@ Abstract
I Introduction
@@
Internet of Things (IoT) environments are characterized by vast amounts of data that are continuously generated by heterogeneous and resource-constrained devices. While this data is valuable for Machine Learning (ML) applications [1], transmitting it to a central server is often impractical due to privacy, bandwidth, and latency constraints. Federated Learning (FL) [2] offers a promising alternative to traditional centralized ML by enabling multiple devices, such as those in the IoT, to collaboratively train shared models without exchanging the training data [3]. In an FL setting, a central server distributes a global model to participating clients that locally compute updates (i.e., gradients or weight differences) using their private dataset. The updates are then sent to the server, which aggregates them to obtain an updated version of the global model that is sent back to the clients. This process is repeated until convergence.
@@
Although FL enhances privacy protection compared to centralized approaches as clients’ data are kept on-site [4], it is still vulnerable to several security and privacy threats [5, 6, 7]. In particular, privacy attacks, which are the main focus of this work, aim to infer sensitive information about the clients, such as training data or their properties, by exploiting the intermediate updates sent to the server [6, 7, 8, 9].
@@
In FL scenarios, server-client communications can be protected through end-to-end encryption, making attacks by external adversaries more difficult. However, this does not prevent an honest-but-curious server from breaking clients’ privacy without being detected, as it follows the regular training protocol while seeking to gain additional information [6, 8]. The server has access to the global model and client updates. Additionally, it is capable of executing computationally intensive reconstruction methods.
@@
Within this threat model, Gradient Inversion Attacks (GIAs) aim to reconstruct the private client data used to compute the model update [8]. These attacks rely on optimization techniques to reconstruct client’s data, with the objective to estimate input values that produce updates closely matching those shared by the victim client [10, 11]. Optimization strategies may also employ some pre-trained generative models like Generative Adversarial Networks (GANs) or diffusion models as strong priors to constrain the optimization process [12, 13, 14]. In the latter case, the attacker must have access to a surrogate dataset to train these models. These datasets should approximate the target client’s data distribution (e.g., if the learning task consists of recognizing faces, then the attacker should have some public photos of other individuals). This assumption is widely accepted in the literature [13, 14, 15, 16, 17, 18, 19, 8].
@@
Figure 1: Comparison of reconstruction results on images when the model update is calculated over 256 images at $224\times 224$ resolution using the VGG-16 model. (a) Reconstruction with CPA [20], (b) Reconstruction with CPA [20] enhanced with GUIDE (ours), (c) Original Images.
GIAs have been demonstrated to be more effective when model updates are computed over small batch sizes or when the data is low-dimensional [10, 21, 11, 13, 17, 14]. However, FL clients usually compute gradients over large batch sizes, and data is high-dimensional (e.g., high-resolution images), making data reconstruction even more challenging. As shown in Figure 1a, when larger batches of high-resolution images are used, existing GIAs often generate noisy approximations of the client’s original private images. This limitation can be overcome if the attacker has access to a surrogate dataset to train tailored ML models aiming to enhance reconstruction quality. To this purpose, recent advancements in generative neural networks, particularly diffusion models [22], hold significant promise to make GIAs more effective.
@@
We implemented GUIDE and conducted extensive evaluations across various state-of-the-art attacks, using different FL algorithms, models, and datasets, while also considering defensive mechanisms. Furthermore, we recognize that commonly used perceptual metrics for assessing GIAs may not be entirely suitable for evaluating privacy leakage in reconstructed images because they do not fully align with human perception [23]. To address this issue, we propose using DreamSim [24], which has recently been established as the state-of-the-art perceptual similarity metric [25]. Experimental results demonstrate that GUIDE significantly enhances the quality of reconstructed data in GIAs, generating images that are both more realistic and semantically closer to the original ones. Specifically, it achieves up to 46% higher perceptual similarity according to the DreamSim metric, with even further gains when certain defensive strategies are applied. Figure 1b visually illustrates how GUIDE enhances the effectiveness of the Cocktail Party Attack (CPA) [20]. These improvements, achieved under the same threat model, directly correlate with increased privacy leakage, even when updates are computed over large batch sizes and high-dimensional data.
@@
  • We demonstrate that the DreamSim [24] metric better assesses the privacy leakage of reconstructed images, offering a more perceptually aligned evaluation of GIAs in FL;
@@
Organization. The remainder of the paper is organized as follows. Section II provides the background and review of related work on GIAs. Section III introduces the considered threat model. Section IV presents GUIDE, whose effectiveness is extensively evaluated in Section V. Finally, Section VI draws our conclusions. Notations and symbols used throughout the paper are clarified in Table I.
@@
TABLE I: Formal notations used in the paper.
II Background & Related Work
@@ II-A Federated Learning
The primary algorithms used in FL are Federated Stochastic Gradient Descent (FedSGD) [2] and Federated Averaging (FedAVG) [2], which mainly differ in how the selected clients $\mathcal{K}\subseteq C$ compute and transmit updates at each iteration to train the global model. In FedSGD [2], each client performs a single local epoch ($E=1$) by sampling a batch from its private dataset ($B\leq N_{c}$) and computing the aggregated gradient $\nabla_{c}=\sum_{i=1}^{B}\nabla\mathcal{L}(x_{c,i},y_{c,i};\theta_{t})$. This computed gradient is then sent to the server, which aggregates the gradients from all selected clients to update the global model as follows:
@@
where $\eta$ is the global learning rate. While FedSGD enables frequent global model updates, it incurs significant communication overhead due to the continuous exchange of gradients [2]. FedAVG [2] improves communication efficiency by allowing clients to perform multiple local training steps ($E>1$) before sharing updates with the server. In each communication round $t$, a client $c\in\mathcal{K}$ initializes its local model as $\theta_{c}^{(t,0)}=\theta_{t}$ and performs $\tau=E\cdot N_{c}/B$ gradient descent steps using a learning rate $\eta$, updating its local model to $\theta_{c}^{(t,\tau)}$. Once local training is completed, the client computes the model update as:
@@
II-B Gradient Inversion Attacks
@@
where $\mathcal{L}_{\text{grad}}$ denotes a loss function that quantifies the discrepancy between the shared update $\nabla$ and the dummy update $\nabla^{\prime}$ computed with respect to the reconstructed input $X^{\prime}$ and labels $Y^{\prime}$. $\mathcal{R}_{\text{aux}}$ is a regularization term that enforces specific constraints on the reconstructed input $X^{\prime}$. The optimization variables $X^{\prime}$ and $Y^{\prime}$ can be initialized either randomly or leveraging prior knowledge relevant to the problem [8].
One of the first attacks is Deep Leakage from Gradients (DLG) [10], which targets small batches and simple architectures under the FedSGD protocol. Subsequent works have demonstrated the effectiveness of similar methods in the presence of more complex FL schemes, such as FedAVG [26, 15]. Progress in this domain has generally involved refining the optimization pipeline with additional priors and constraints. For instance, InvGrad [11] uses a combination of cosine distance and total variation regularization, thereby enabling the inversion of deeper networks and larger batches. GradInversion [21] further exploits Batch Normalization (BN) statistics to guide reconstructions, aligning the activations produced by candidate inputs with the global BN statistics stored in the model. Meanwhile, some works isolate label reconstruction as a separate challenge, which can be addressed analytically (e.g., assuming unique labels per batch). In another direction, CPA [20] casts gradient inversion in Fully Connected (FC) layers as a blind source separation problem, first recovering private embeddings by unmixing the FC-layer gradients via techniques akin to independent component analysis, and then reconstructing images through an auxiliary optimization.
Some other attacks [12, 13, 16, 27] leverage pre-trained generative models to reduce the complexity of searching directly in the high-dimensional image space. Concretely, these approaches restrict the optimization process to a latent code $z^{\prime}\in\mathcal{Z}$, under the assumption of having a generator $G\colon\mathcal{Z}\to\mathcal{X}$ that maps latent vectors to candidate images. The server-side updates $\nabla^{\prime}$ are then computed by substituting $G(z^{\prime})$ into the client’s model in place of the actual private inputs. Formally, one solves:
@@
Recently, [14] have proposed a novel GIA based on the latent-space optimization principle (Eqn. 7), employing a diffusion model as the pre-trained generative model instead of a GAN. It is worth noting that the diffusion model is used as a generative prior, guiding the iterative synthesis process from pure noise by incorporating the shared gradient as a directive condition at each sampling step. This significantly differs from our methodology that leverages the diffusion model as a denoising tool specialized for the underlying GIA. Additionally, [14] was validated on model updates computed from a single image (a batch size of one), which is an unrealistic setting for practical FL deployments.
II-C Augmented Optimization Attacks
Other strategies seek to improve GIAs by incorporating image restoration modules, particularly when defensive measures based on gradient obfuscation are in place [8]. Highly-Compressed Gradient Leakage Attack (HC-GLA) [17] combines standard gradient-based reconstruction steps with a specialized Convolutional Neural Network (CNN) for noise suppression. This method differs from state-of-the-art image restoration models, as it is a CNN trained on a limited dataset of only a few thousand images, which restricts its overall potential and effectiveness in comparison to more advanced approaches. Moreover, it is primarily evaluated on single-image batches and may struggle to scale or generalize to diverse data distributions or larger batch sizes.
Reconstructing from Obfuscated Gradient (ROG) [15] attack introduces a two-step GIA designed to enhance both convergence and visual fidelity. First, the attack operates in a reduced-dimensional space by encoding the input data, simplifying the optimization objective, and accelerating convergence. Once a suitable representation is found, a GAN is used to restore the optimized embedding into the original high-dimensional image space. Unlike GUIDE, this GAN is not specifically trained on the intermediate noisy reconstructions produced by the initial attack step. This limits its adaptability and results in a significant performance drop when targeting deeper neural architectures.
@@ II-C Augmented Optimization Attacks
III Threat Model
In this section, we provide a comprehensive definition of the threat model adopted in our work, specifying the nature of the adversary, along with its objectives, knowledge, and capabilities. In line with the literature on GIAs [8, 9], we consider the central server as the adversary.
Adversary Goal. The main goal of the central server is to break clients’ privacy by gaining information about their private data used for the training.
@@ III Threat Model
In some prior works, the use of a surrogate dataset is not explicitly stated; however, this assumption is implicitly made through the application of pre-trained generative models. Specifically, this threat model closely aligns with one of the threat models defined in [8], followed by several works in literature [13, 14, 15, 16, 17, 18, 19].
Adversary Capability. The central server in FL could be either active malicious or honest-but-curious [8]. An active malicious server intentionally deviates from the protocol to compromise client privacy by manipulating the global model to facilitate data reconstruction. In contrast, we adopt the honest-but-curious server, a more realistic threat model in FL deployments. This reflects scenarios where the server faithfully follows the FL protocol, but passively attempts to infer private information from clients’ updates [8]. It is worth noting that, under this threat model, clients are unable to detect the server’s malicious behavior.
IV GUIDE
@@
IV-A Intuition behind GUIDE
Several works have experimentally shown the effectiveness of GIAs to reconstruct an approximated version of the client’s input, especially when additional priors constrain the optimization procedure [8]. However, there is no theoretical guarantee that these reconstructions correspond to the data used by the client during the local training. Indeed, GIAs based on gradient-matching optimization procedures (e.g., those based on Equations 6 and 7) are inherently ill-posed mathematical problems since they potentially admit several solutions consistent with the observed client’s gradient [28].
The reconstructed inputs obtained through GIAs often exhibit noisy artifacts, which can be attributed to the fact that the optimization may place the reconstructed points in lower-density regions of the data manifold, where the original client data is less likely to reside. Thus, the recovered points may satisfy the observed gradients but still deviate from the true data distribution. Given these considerations, we leverage generative image enhancement models trained with a surrogate dataset to move approximated input reconstructions into regions of the data space that are more likely to contain the original client inputs [29]. This “guides” the optimization process toward high-density areas of the learned distribution (hence the name of the method), reducing ambiguity in the reconstruction and improving the overall fidelity of the recovered samples.
@@ IV-A Intuition behind GUIDE
IV-B GUIDE Setup